OpenStudy (acespeedfighter):

If the side lengths of a cube are 14 feet, what is the correct way to write the expression to represent the volume of the cube in exponential form? (2 points) 143 314 14 ⋅ 3 14 ⋅ 14 ⋅ 14 2. (06.01) Simplify the expression 43 + 2(3 − 2). (3 points) 14 16 66 68 3. (06.04) What is the value of 4 + 5x when x = 3? (3 points) 11 12 17 19

OpenStudy (acespeedfighter):

14x3 right?

OpenStudy (acespeedfighter):

for numbero uno

MARC (marc_d):

cube-->all sides are equal length

OpenStudy (acespeedfighter):

for D 14x14x14

MARC (marc_d):

formula for volume of cube is lengthx widthx height

OpenStudy (acespeedfighter):

So its D :D

MARC (marc_d):

\[=14\times14\times14\]

MARC (marc_d):

yess

OpenStudy (acespeedfighter):

got emm

MARC (marc_d):

okay..second question..

OpenStudy (acespeedfighter):

I dont under stand what the question is asking

OpenStudy (yanasidlinskiy):

Just to clarify they're parentheses not brackets. Those are two different stories

OpenStudy (acespeedfighter):

Oh so divide?

OpenStudy (acespeedfighter):

I mean SUbstrat

MARC (marc_d):

thanks for the reminder @YanaSidlinskiy\[43+2(3-2)\]\[43+2(1)\]\[43+2\]

OpenStudy (acespeedfighter):

Best Speller^^^

OpenStudy (acespeedfighter):

hmmmm

OpenStudy (acespeedfighter):

what am i going to do divide 43 by 2?

MARC (marc_d):

no,u add them together..

OpenStudy (acespeedfighter):

45...

MARC (marc_d):

yep,that's right.

OpenStudy (acespeedfighter):

but tahts not a choice

MARC (marc_d):

hmm

MARC (marc_d):

the working should be correct..

MARC (marc_d):

\[=43+2(3-2)\]\[=43+6-4\]\[=43+2\]\[=45\]it should be like this

OpenStudy (acespeedfighter):

look at the answers

MARC (marc_d):

yea,i know that.. xD

OpenStudy (acespeedfighter):

XD

OpenStudy (acespeedfighter):

Simplify the expression 43 + 2(3 − 2). (3 points) 14 16 66 68 Nope

MARC (marc_d):

@YanaSidlinskiy what do u think?

OpenStudy (acespeedfighter):

@pooja195

OpenStudy (yanasidlinskiy):

I got 45 as well. There may not be an answer choice for this

OpenStudy (acespeedfighter):

those are the only choices

MARC (marc_d):

agree with u @YanaSidlinskiy i guess there is some error..

OpenStudy (welshfella):

do the part thats in the parentheses first (the 3 - 2) then multiply by finally add the 43

OpenStudy (welshfella):

* multiply by 2

OpenStudy (acespeedfighter):

so 45 but the answers are this

OpenStudy (acespeedfighter):

Simplify the expression 43 + 2(3 − 2). (3 points) 14 16 66 68

OpenStudy (acespeedfighter):

14, 16, 66, 68

pooja195 (pooja195):

One question per post man. And there's people trying to help you please cooperate with them rather than ignoring them.

OpenStudy (welshfella):

as its written the answer is definietely 45 Theres a mistake somewhere.

OpenStudy (yanasidlinskiy):

Choose your best option is all I could say

pooja195 (pooja195):

I got 45... did you exclude something from the question?

OpenStudy (acespeedfighter):

Look at that

pooja195 (pooja195):

There's a mistake in how you wrote it.

OpenStudy (acespeedfighter):

look at the screen shot

pooja195 (pooja195):

Its 4^3 not 43.

OpenStudy (acespeedfighter):

XD there we go

pooja195 (pooja195):

\[\huge~\rm~\bf~4^3 + 2(3 − 2)\]

OpenStudy (acespeedfighter):

66 1 so add these and get 67

OpenStudy (acespeedfighter):

wait

OpenStudy (acespeedfighter):

is it 68?

OpenStudy (yanasidlinskiy):

Then you posted the question wrong itself. The exponent makes a WHOLE LOT of difference.

pooja195 (pooja195):

Its not 68 4^3=?

OpenStudy (acespeedfighter):

oh 66

pooja195 (pooja195):

Yep good job

OpenStudy (acespeedfighter):

Thx sorry for the error